### v2

@@ 5.2 Linear regression sampling
Another key problem type for which it is possible to transform data into a rival good is sampling problems with a quantum communication advantage. In this type of problem, Alice sends Bob a message, which Bob uses to sample from a target distribution with high accuracy. The essence of the construction is that the quantum communication advantage allows Alice to reveal only a tiny fraction of the original data while allowing Bob to solve the problem, and the method by which he solves it destroys the data that was sent, not allowing it to be reused to generate more samples. We consider here a sampling variant of linear regression introduced by Montanaro et al. [39]:
Problem 3 (Linear Regression Sampling [39] ($\mathrm{LRS}_{N}$)).
@@
Problem 4 (Multiple Linear Regression Sampling ($\mathrm{MLRS}_{N,m}$)).
@@
Note that solving the above problem with some inaccuracy $\eta$ corresponds to sampling from some distribution with total variation error at most $\eta$ with respect to $\mathcal{P}_{k}$. In order to consider the communication complexity of these problems, we must first discretize the inputs so that they have finite size. We thus assume all real number are specified to $\log N$ bits of precision. We then have
Lemma 9 ([39]).
@@
5.3 Two-outcome observable estimation
In the previous examples, we considered problems that exhibit exponential quantum communication advantages. It is natural to ask if such an advantage implies consumability in some generic sense. We will see that this is not the case when Bob’s task is a decision problem.
@@ 5.3 Two-outcome observable estimation
Problem 5 (Vector In Subspace ($\mathrm{VS}_{N,\theta}$) [32]).
@@
It is known that Vector in Subspace exhibits an exponential advantage in quantum communication with respect to randomized classical communication complexity [44]. Consider the following generalization:
Problem 6 (Vector In Multiple Subspaces ($\mathrm{VMS}_{N,\theta,m}$)).
@@
Theorem 1 (Shadow Tomography [2] solved with Threshold Search [14]).
@@
$\mathrm{VMS}_{N,\theta,m}$ is a problem of estimating $m$ expectation values up to some constant error (due to the constraint on $\theta$) on a target state. If polynomial error is required, it is known that $\Omega(N)$ qubits of communication may be needed, and hence quantum communication is essentially equivalent to classical communication (from e.g. lower bounds on estimating inner products [18]). Allowing constant error, Theorem 1 says that $\mathrm{polylog}(m)$ qubits of communication suffice to solve the problem. This directly implies that, at least if Alice sends multiple copies of her state, a lower bound analogous to Lemma 10 is impossible, as Bob does not require a number of qubits polynomial in $m$. This shows that an exponential communication advantage is not a sufficient condition for quantum data to behave like a rival good.
Given that multiple entangled copies of a quantum state are known to be a more powerful resource than single copies [27], it would also be interesting to consider a setting where Alice sends only single copies of her data states. One way to do this is by introducing assumptions about the computation Bob is allowed to perform with his message. It may be possible to remove this assumption by utilizing certified deletion [12]. While requiring additional encryption, this could enable Alice to only send a copy of her state after receiving a certificate that Bob has deleted the previous copy, ensuring that multi-copy measurements cannot be performed.
@@ 5.4 Multiple Bobs: A communication arms race
The above picture changes when more than two parties are involved. Consider a setting where Alice has a vector which she can encode in a quantum state |$\ket{x}$ and each of $m$ Bobs has an observable $O_{i}$, Alice is only willing to send the Bobs copies of |$\ket{x}$ (when using quantum communication), and the Bobs cannot (i) store multiple copies of |$\ket{x}$ or (ii) communicate quantum states between them, this is equivalent to the setting of learning without quantum memory that is studied in [17]. More precisely, this is a setting where each Bob can perform a POVM on a single copy of |$\ket{x}$ only, and exchange classical messages which correspond to the classical memory used in this setting. In contrast, the setting of learning with quantum memory (as per [17]) is one where the Bobs are allowed quantum communication (but still can measure only a single copy of |$\ket{x}$ each), with the content of the quantum communication channel corresponding to the quantum memory. In both cases, Alice’s messages correspond to samples of a quantum state (unknown to Bob) as is standard in learning problems. While the results of [17] apply to samples of a mixed state described by a density matrix $\rho$, they also apply to a purification of $\rho$ in a larger space. This will not affect the scaling with $m$ which is the main object of interest for our purposes.
@@
When only classical communication is used between Alice and the Bobs, an optimal lower bound of $\Omega(\sqrt{N})$ for estimating the expectation value of a single two-outcome observable with constant probability is applicable [25]. Lemma 1 of that paper also provides a matching upper bound in the $m$-observable case (up to logarithmic factors). Namely, estimating $m$ expectation values of unit norm observables to constant error can be done with probability $2/3$ by sending $\tilde{O}(\log(m)\sqrt{N})$ bits from Alice to Bob (where $\tilde{O}$ hides $\mathrm{polylog}(N)$ factors). Alice requires no knowledge of the observables themselves. This protocol is based on sending $O(\log(m))$ random stabilizer sketches of Alice’s input state |$\ket{x}$. Each sketch involves Alice drawing a Clifford unitary $C$ from a uniform distribution over the Clifford group $\mathcal{C}_{n}$ ($n=\log N)$, and computing $\bigl{\langle}0^{\otimes(n-k)}z\bigr{|}C\bigl{|}x\bigr{\rangle}$ for all $z\in\ket{0,1}^{k}$ for $2^{k}=\tilde{O}(\sqrt{N})$. Alice generates $O(\log(m))$ i.i.d. sketches in this way and sends both the measurement results and a description of the Clifford unitaries to the Bobs. Each Clifford unitary is defined by specifying $O(n^{2})$ one or two-qubit gates from a small set, and thus has an efficient classical description.
If Alice instead sends copies of her input encoded in the amplitudes of a quantum state |$\ket{x}$ to the Bobs, but we allow classical communication only between the Bobs, and restrict the Bobs to performing single-copy measurements, the number of samples of |$\ket{x}$ required is linear in $m$ [17]:
Theorem 2 (Corollary 5.7, [17]).
@@
Classical A $\rightarrow$ Bs, Classical Bs $\leftrightarrow$ Bs $\tilde{\Theta}(N^{1/2})$ [25]
Quantum A $\rightarrow$ Bs, Classical Bs $\leftrightarrow$ Bs $\tilde{\Theta}(m\log N)$ [28]
Quantum A $\rightarrow$ Bs, Quantum Bs $\leftrightarrow$ Bs $O((\log(m)\log(N))^{2})$ [14]
Table 1: A communication arms race in estimating expectation values of two-outcome observables to within constant error: Data behaves as a consumable resource if Alice is only willing to send quantum states encoding her data, while the Bobs can only communicate classically. This ceases to be the case if only classical communication is used, or if the Bobs can communicate quantum states. $\tilde{\Theta}$ hides factors of $\log m$.
Note that this is worst-case over |$\ket{x}$ (if |$\ket{x}$ was uniformly random Bobs could just guess 00). Note also that the $O_{i}$ are chosen so that classical shadows do not help (for the operators in question the Hilbert-Schmidt norm is $||O^{2}_{i}||=N$, which is roughly equivalent to the shadow norm that sets the sample complexity of classical shadows [28]). A matching upper bound (up to $\log(m)$ factors, as long as $m<N$) is obtained by the straightforward approach in which Alice sends each Bob $O(1/\varepsilon_{2})$ copies of her state.
When the Bobs are allowed to use quantum communication, we are essentially back to the two-party version of the problem, since they can jointly use shadow tomography [2, 14] to estimate all the expectation values using a logarithmic number of copies of |$\ket{x}$. These results are summarized in Table 1.
@@ 6.1 Data as an economic resource in production theory
Production theory [34] is one of the principal frameworks for the quantitative study of economic systems. A fundamental object of interest within this framework is the production function $F:\mathbb{R}_{+}^{M}\rightarrow\mathbb{R}_{+}$ that quantifies in some form the output of an economic agent, for example the goods produced by a firm. The inputs to $F$ denote the resources required to produce said goods, such as labor, capital and raw materials. For conventional goods of this form, which cannot be replicated at zero cost (and are referred to as rival goods), it is known that the production function is typically a degree $1$ homogeneous function of its inputs (at least locally when restricted to some set $S$):
@@
for all $i\in[M]$, then the output of the firm suffices exactly to purchase all the resources required, and there is no surplus profit. This is known as competitive equilibrium, which maximizes social welfare in the sense that the price of each good is commensurate to its usefulness in increasing the total output [7, 19].
While it has long been understood at a qualitative level that data is an inherently different resource than the ones considered above due to the ability to copy it for free [47, 8], the quantitative form of this statement was realized decades later by the seminal work of Romer [45]. If we include data $y$ as an input into the production function, we instead have
@@
Due to this inequality, it is impossible to set prices according to Eq. 6.3. If this were done for all inputs including data, the total output would be insufficient to pay for all the required resources. As a result, markets involving data must be inherently inefficient in the sense that one must underpay for some resource, or must include some external mechanism to enforce adequate compensation for resources that can be freely replicated. Mechanisms such as patent law or subsidies that incentivize innovation are all examples of this. Other examples are afforded by the trusted third parties that are introduced in proposals for data markets and handle the data in lieu of the data buyers themselves [5]. In the context of strategic games that model data selling, the ability to copy data is also manifest in the payoff for the data seller being independent of the number of buyers, unless a mechanism is put in place by which the data buyers all agree to pay in advance for their data [40].
@@ 6.2 A posted price data auction with consumable data
We would like to identify more concretely the economic consequences of the consumable nature of quantum data. We consider a formulation naturally related to auction theory [33, 46]. Alice’s action space $A_{A}=\mathbb{R}_{+}$ is the set of prices she charges for a single bit or qubit of her input. Once Alice fixes a price $p$, Bob is free to purchase as many bits/qubits as he wants. Bob’s action space is thus $A_{B}=\mathbb{N}$, and we denote the number he purchases by $b$. This is known as a posted price auction with only a single bidder and multiple items (or a particularly simple combinatorial auction). Assume the number of samples $m$ takes values in $[\overline{m}]$ and Alice has no knowledge of it (say she holds a uniform prior). We also assume the matrices $B_{i}$ are chosen in a worst-case fashion (in order for our communication lower bounds to be applicable).
@@
In the classical case, we know the problem is nonconsumable from Lemma 2. This implies that for $m=1$, there is a message of length $\kappa$ independent of $m$ which Alice can send, which Bob can then re-use to produce say $\rho m$ samples with some constant probability, for some $\rho\leq 1$.
@@
7 Discussion
Being a preliminary investigation into the possibility of using quantum networks in this manner, our results do not immediately apply to problems with clear economic value. If this were the case, it could enable novel types of data markets and incentive structures for the production of data. It is worth noting however that our results for the linear regression sampling problem apply also to a related problem in which Bob obtains a state that encodes the solution to a linear system rather than a classical sample. Such states are known to be strictly more powerful resources than classical samples [6], and could potentially be useful in learning tasks such as updating the value of a linear estimator with new data (which is typically achieved with the recursive least squares algorithm).
The form of the quantum communication lower bound that indicates the rival behavior of quantum data is reminiscent of a direct sum theorem. Direct sum theorems demonstrate that the complexity of solving $m$ independent instances of certain problems scales linearly with $m$. They have been studied extensively in both the classical [48, 11, 37] and quantum [49, 29] setting. These results are not directly applicable since in our setting the inputs to Alice are not independent. Thus, this work motivates an asymmetric direct sum result for classes of communication relations.
In restricting access to data that is used for computation, the setting we consider bears some resemblance to that of differential privacy [20, 4]. In differential privacy, a query is promised not to reveal too much about individual datapoints. This is typically achieved classically by adding noise to data, while we achieve a similar capability in spirit by using a noiseless encoding into quantum states.
7.1 Open questions
@@
A remark regarding question 3 is that for a random relation, it is actually impossible to get a constant probability of success for superconstant number of instances because we do not have amplification of success probability. This is why we ask for explicit examples.
Our results are unconditional but restricted to specific problems and one-way communication. By making additional assumptions or utilizing the strategic nature of the problems we consider, it may be possible to extend the class of problems that enable consumable data. Some possible directions are outlined below
Computational assumptions: A setting we have not yet considered, that is touched upon by the task of shadow tomography, is one where the computational power of Bob is restricted. It has been noted that general shadow tomography procedures are expected to scale polynomially with the dimension of the Hilbert space of $\rho$ or the trial state $\rho_{T}$. If Bob is restricted to polylog computational time, then the creation of the clonable hypothesis state may become impossible. This is analogous to the effect in cryptographic no-cloning theorems on pseudorandom quantum states [30], where even when sample efficient cloning is possible, no computationally efficient scheme can be used to clone the states of interest. In this context, it is worth noting that learning of certain states that have efficient descriptions, such as pseudorandom states [54], is known to be computationally hard. The addition of computational restrictions on Bob hence potentially widens the class of consumable data tasks, but requires moving beyond a communication complexity model that permits unbounded computation.
Communicating mixed states: Moreover, in the above schemes we have focused on cases where the data is provided as pure states and when the only advantage examined otherwise is communication complexity. When data is transmitted as a pure state, it is known that classical shadows have the potential to remove strict communication advantages when the circuits are too simple, but reconstruction can still be challenging computationally based on arguments from quantum pseudorandom states [54]. An interesting direction could be to enhance some of the features of the example problems here by providing the data as a mixed state, such that it is more difficult for an adversary to learn under certain assumptions such as lack of a substantial quantum memory. We leave these directions for future work.

### v3

@@ 5.2 Linear regression sampling
Another key problem type for which it is possible to transform data into a rival good is sampling problems with a quantum communication advantage. In this type of problem, Alice sends Bob a message, which Bob uses to sample from a target distribution with high accuracy. The essence of the construction is that the quantum communication advantage allows Alice to reveal only a tiny fraction of the original data while allowing Bob to solve the problem, and the method by which he solves it destroys the data that was sent, not allowing it to be reused to generate more samples. We consider here a sampling variant of linear regression introduced by Montanaro et al. [MS24]:
Problem 4 (Linear Regression Sampling [MS24] ($\mathrm{LRS}_{N}$)).
@@
Problem 5 (Multiple Linear Regression Sampling ($\mathrm{MLRS}_{N,m}$)).
@@
Note that solving the above problem with some inaccuracy $\eta$ corresponds to sampling from some distribution with total variation error at most $\eta$ with respect to $\mathcal{P}_{k}$. In order to consider the communication complexity of these problems, we must first discretize the inputs so that they have finite size. We thus assume all real number are specified to $\log N$ bits of precision. We then have the following lemma for classical protocols.
Lemma 10 ([MS24]).
@@
Meanwhile for quantum protocols, we have a linear dependence on $m$.
@@
5.3 Decision problems
In the previous examples, we considered problems that exhibit an exponential quantum communication advantage. It is natural to ask if such an advantage implies consumability in some generic sense. We will see that this is not the case when Bob’s task is a decision problem.
@@ 5.3 Decision problems
Problem 6 (Vector In Subspace ($\mathrm{VS}_{N,\theta}$) [Kre95]).
@@
It is known that Vector in Subspace exhibits an exponential advantage in quantum communication with respect to randomized classical communication complexity [RK11]. Consider the following generalization:
Problem 7 (Vector In Multiple Subspaces ($\mathrm{VMS}_{N,\theta,m}$)).
@@
Theorem 1 (Shadow Tomography [Aar18] solved with Threshold Search [BO21]).
@@
$\mathrm{VMS}_{N,\theta,m}$ is a problem of estimating $m$ expectation values up to some constant error (due to the constraint on $\theta$) on a target state. If polynomial error is required, it is known that $\Omega(N)$ qubits of communication may be needed, and hence quantum communication is essentially equivalent to classical communication (from e.g. lower bounds on estimating inner products [CvDNT99]). Allowing constant error, Theorem 1 says that $\mathrm{polylog}(m)$ qubits of communication suffice to solve the problem. This directly implies that, at least if Alice sends multiple copies of her state, a lower bound analogous to Lemma 11 is impossible, as Bob does not require a number of qubits polynomial in $m$. This shows that an exponential communication advantage is not a sufficient condition for quantum data to behave like a rival good.
Given that multiple entangled copies of a quantum state are known to be a more powerful resource than single copies [HBC+22], it would also be interesting to consider a setting where Alice sends only single copies of her data states. One way to do this is by introducing assumptions about the computation Bob is allowed to perform with his message. It may be possible to remove this assumption by utilizing certified deletion [BI20]. While requiring additional encryption, this could enable Alice to only send a copy of her state after receiving a certificate that Bob has deleted the previous copy, ensuring that multi-copy measurements cannot be performed.
@@ 5.4 Multiple Bobs: A communication arms race
The above picture changes when more than two parties are involved. Consider a setting where Alice has a vector which she can encode in a quantum state |$\ket{x}$ and each of $m$ Bobs has an observable $O_{i}$, Alice is only willing to send the Bobs copies of |$\ket{x}$ (when using quantum communication), and the Bobs cannot (i) store multiple copies of |$\ket{x}$ or (ii) communicate quantum states between them, this is equivalent to the setting of learning without quantum memory that is studied in [CCHL22]. More precisely, this is a setting where each Bob can perform a POVM on a single copy of |$\ket{x}$ only, and exchange classical messages which correspond to the classical memory used in this setting. In contrast, the setting of learning with quantum memory (as per [CCHL22]) is one where the Bobs are allowed quantum communication (but still can measure only a single copy of |$\ket{x}$ each), with the content of the quantum communication channel corresponding to the quantum memory. In both cases, Alice’s messages correspond to samples of a quantum state (unknown to Bob) as is standard in learning problems. While the results of [CCHL22] apply to samples of a mixed state described by a density matrix $\rho$, they also apply to a purification of $\rho$ in a larger space. This will not affect the scaling with $m$ which is the main object of interest for our purposes.
@@
When only classical communication is used between Alice and the Bobs, an optimal lower bound of $\Omega(\sqrt{N})$ for estimating the expectation value of a single two-outcome observable with constant probability is applicable [GS19]. Lemma 1 of that paper also provides a matching upper bound in the $m$-observable case (up to logarithmic factors). Namely, estimating $m$ expectation values of unit norm observables to constant error can be done with probability $2/3$ by sending $\tilde{O}(\log(m)\sqrt{N})$ bits from Alice to Bob (where $\tilde{O}$ hides $\mathrm{polylog}(N)$ factors). Alice requires no knowledge of the observables themselves. This protocol is based on sending $O(\log(m))$ random stabilizer sketches of Alice’s input state |$\ket{x}$. Each sketch involves Alice drawing a Clifford unitary $C$ from a uniform distribution over the Clifford group $\mathcal{C}_{n}$ ($n=\log N)$, and computing $\bigl{\langle}0^{\otimes(n-k)}z\bigr{|}C\bigl{|}x\bigr{\rangle}$ for all $z\in\ket{0,1}^{k}$ for $2^{k}=\tilde{O}(\sqrt{N})$. Alice generates $O(\log(m))$ i.i.d. sketches in this way and sends both the measurement results and a description of the Clifford unitaries to the Bobs. Each Clifford unitary is defined by specifying $O(n^{2})$ one or two-qubit gates from a small set, and thus has an efficient classical description.
If Alice instead sends copies of her input encoded in the amplitudes of a quantum state |$\ket{x}$ to the Bobs, but we allow classical communication only between the Bobs, and restrict the Bobs to performing single-copy measurements, the number of samples of |$\ket{x}$ required is linear in $m$ [CCHL22]:
Theorem 2 (Corollary 5.7, [CCHL22]).
@@
Classical A $\rightarrow$ Bs, Classical Bs $\leftrightarrow$ Bs $\tilde{\Theta}(N^{1/2})$ [GS19]
Quantum A $\rightarrow$ Bs, Classical Bs $\leftrightarrow$ Bs $\tilde{\Theta}(m\log N)$ [HKP20]
Quantum A $\rightarrow$ Bs, Quantum Bs $\leftrightarrow$ Bs $O((\log(m)\log(N))^{2})$ [BO21]
Table 1: A communication arms race in estimating expectation values of two-outcome observables to within constant error: Data behaves as a consumable resource if Alice is only willing to send quantum states encoding her data, while the Bobs can only communicate classically. This ceases to be the case if only classical communication is used, or if the Bobs can communicate quantum states. $\tilde{\Theta}(\cdot)$ hides factors of $\log m$.
Note that this is worst-case over |$\ket{x}$ (if |$\ket{x}$ was uniformly random Bobs could just guess 00). Note also that the $O_{i}$ are chosen so that classical shadows do not help (for the operators in question the Hilbert-Schmidt norm is $||O^{2}_{i}||=N$, which is roughly equivalent to the shadow norm that sets the sample complexity of classical shadows [HKP20]). A matching upper bound (up to $\log(m)$ factors, as long as $m<N$) is obtained by the straightforward approach in which Alice sends each Bob $O(1/\varepsilon_{2})$ copies of her state.
When the Bobs are allowed to use quantum communication, we are essentially back to the two-party version of the problem, since they can jointly use shadow tomography [Aar18, BO21] to estimate all the expectation values using a logarithmic number of copies of |$\ket{x}$. These results are summarized in Table 1.
@@ 6 Economic implications of consumable data
We now illustrate the economic implications of the complexity-theoretic asymmetric direct sum property. We give some background on the effect of the ability to copy classical data on economic models, and, more concretely, design an auction where the buyer is forced to purchase $m$ copies of the data if they want to use it $m$ times, increasing the payoff of the data seller compared to an analogous situation when selling classical data.
@@ 6.1 Data as an economic resource in production theory
Production theory [KS95] is one of the principal frameworks for the quantitative study of economic systems. A fundamental object of interest within this framework is the production function $F:\mathbb{R}_{+}^{M}\rightarrow\mathbb{R}_{+}$ that quantifies in some form the output of an economic agent, for example the goods produced by a firm. The inputs to $F$ denote the resources required to produce said goods, such as labor, capital and raw materials. For conventional goods of this form, which cannot be replicated at zero cost (and are referred to as rival goods), it is known that the production function is typically a degree $1$ homogeneous function of its inputs (at least locally when restricted to some set $S$):
@@
for all $i\in[M]$, then the output of the firm suffices exactly to purchase all the resources required, and there is no surplus profit. This is known as competitive equilibrium, which maximizes social welfare in the sense that the price of each good is commensurate to its usefulness in increasing the total output [Arr51, Deb59].
While it has long been understood at a qualitative level that data is an inherently different resource than the ones considered above due to the ability to copy it for free [Sch42, Arr62], the quantitative form of this statement was realized decades later by the seminal work of Romer [Rom90]. If we include data $y$ as an input into the production function, we instead have
@@
Due to this inequality, it is impossible to set prices according to Eq. 6.3. If this were done for all inputs including data, the total output would be insufficient to pay for all the required resources. As a result, markets involving data must be inherently inefficient in the sense that one must underpay for some resource, or must include some external mechanism to enforce adequate compensation for resources that can be freely replicated. Mechanisms such as patent law or subsidies that incentivize innovation are all examples of this. Other examples are afforded by the trusted third parties that are introduced in proposals for data markets and handle the data in lieu of the data buyers themselves [ADS19]. In the context of strategic games that model data selling, the ability to copy data is also manifest in the payoff for the data seller being independent of the number of buyers, unless a mechanism is put in place by which the data buyers all agree to pay in advance for their data [NACZL20].
@@ 6.2 A posted price data auction with consumable data
We would like to identify more concretely the economic consequences of the consumable nature of quantum data. We consider a formulation naturally related to auction theory [Kri09, Rou16]. Alice’s action space $A_{A}=\mathbb{R}_{+}$ is the set of prices she charges for a single bit or qubit of her input. Once Alice fixes a price $p$, Bob is free to purchase as many bits/qubits as he wants. Bob’s action space is thus $A_{B}=\mathbb{N}$, and we denote the number he purchases by $b$. This is known as a posted price auction with only a single bidder and multiple items (or a particularly simple combinatorial auction). Assume the number of samples $m$ takes values in $[\overline{m}]$ and Alice has no knowledge of it (say she holds a uniform prior). We also assume the matrices $B_{i}$ are chosen in a worst-case fashion (in order for our communication lower bounds to be applicable).
@@
In the classical case, we know the problem is non-consumable from Lemma 2. This implies that for $m=1$, there is a message of length $\kappa$ independent of $m$ which Alice can send, which Bob can then re-use to produce say $\rho m$ samples with some constant probability, for some $\rho\leq 1$.
@@
7 Open questions
@@
  4. 4.
    We considered the scenario where the graph of required computations is a star graph, where one input must be paired with every other input. You can more generally consider the asymmetric direct sum for any bipartite graph. It is easy to see that the proof of Lemma 8 generalizes to say that the quantum communication complexity scales with the sum of the square-roots of the degrees of the vertices on the left. Can this be improved?
A remark regarding question 3 is that for a random relation, it is actually impossible to get a constant probability of success for every Bob for superconstant number of Bobs because we do not have amplification of success probability. This is why we ask for explicit examples. The lower bound technique would also need to avoid the trivial upper bound if we only require a constant fraction of Bobs to succeed (Lemma 4).
Our results are unconditional but restricted to specific problems and one-way communication. By making additional assumptions or utilizing the strategic nature of the problems we consider, it may be possible to extend the class of problems that enable consumable data. One possible direction is outlined below. Moreover, in Appendix B we explore the broader implications of our results beyond the scope of complexity theory.
Computational assumptions.
A setting we have not yet considered, that is touched upon by the task of shadow tomography, is one where the computational power of Bob is restricted. It has been noted that general shadow tomography procedures are expected to scale polynomially with the dimension of the Hilbert space of $\rho$ or the trial state $\rho_{T}$. If Bob is restricted to polylog computational time, then the creation of the clonable hypothesis state may become impossible. This is analogous to the effect in cryptographic no-cloning theorems on pseudorandom quantum states [JLS18], where even when sample efficient cloning is possible, no computationally efficient scheme can be used to clone the states of interest. In this context, it is worth noting that learning of certain states that have efficient descriptions, such as pseudorandom states [ZLK+23], is known to be computationally hard. The addition of computational restrictions on Bob hence potentially widens the class of consumable data tasks, but requires moving beyond a communication complexity model that permits unbounded computation.